The far-zone electric field of an antenna can be approximated byP.11-add3 The far-zone electric field of an antenna can be approximated by le jkr Je cos? 0, 0sos E. = { r otherwise E = 0. F
Answer #1


Solution . SPE . - cose Osos TV ED - otherwise ( PERE) Ep=0 (a) The Halp The Half power Beam Width in the YZ& Vz-plones Jo.70 = (Bilalyz = 33 & 66° is e eco 020 6 The directivity Do 0 ID= Umax Prad 4 Umax = ? Prad = ? nolulu dr W JT D. Prad - 2 TTV co-jkr Sarg 4x e costo (d. Savg = 1x leikr cosorar n eo-coset isinn I count sina avg - Incosto COSO il p V r imo!! ! ar vom Pro30 plinitril heregge le Pacliation Intensity Su음 - Say x2 so v mask Umax = + (%) max = 1-0 equcetion 2&3 Substitute in equati③ Aproximate directivity Dousing equation Do= 0.78 y 4T (B/2 ) xz (B/2) y z u Byz) x2 = XZ 33 x 1/180 s ets @ ya)gz - 66x T1/